# TRecViT: A Recurrent Video Transformer

### v1

@@ Abstract
We propose a novel block for video modelling. It relies on a time–space–channel factorisation with dedicated blocks for each dimension: gated linear recurrent units (LRUs) perform information mixing over time, self-attention layers perform mixing over space, and MLPs over channels. The resulting architecture TRecViT performs well on sparse and dense tasks, trained in supervised or self-supervised regimes. Notably, our model is causal and outperforms or is on par with a pure attention model ViViT-L on large scale video datasets (SSv2, Kinetics400), while having 3×3$\times$3 × less parameters, 12×12$\times$12 × smaller memory footprint, and 5×5$\times$5 × lower FLOPs count. Code and checkpoints are available online.111 https://github.com/google-deepmind/trecvit
Figure 1: Left: TRecViT architecture. Each video frame is divided into non-overlapping patches that are linearly projected into a token embedding space. We then add a learnt spatial positional encoding. The tokens are passed through gated linear recurrent units (LRUs) that share parameters across space. The outputs of the recurrent blocks are then processed by a ViT block. The recurrent operation followed by ViT is repeated N times. Right: TRecViT block. The input is a batch of videos, each frame with N tokens. We apply recurrent units over temporal tubes to integrate information over time, and self-attention and MLP across tokens within each frame. Note that the recurrent units share parameters, but the information is not mixed across temporal tubes. Similarly, the ViT blocks share parameters, but the information is not mixed across frames.
@@ 1 Introduction
Video understanding requires low-level scene understanding (e.g. how objects move) and high-level reasoning (e.g. causal relations between events) over a signal that is high-dimensional, can be noisy, and contains high correlations and redundancies in both spatial and temporal dimensions. Efficient video modelling needs high-capacity models that can represent the sheer diversity and richness of real-world videos, while having reasonable compute and memory footprint both at training and during inference time. Convolutional neural networks [8, 14] have been a successful family of models for video, but their scaling capabilities (in both data and parameters) are limited due to their inductive biases (locality, invariance). Recurrent neural networks, e.g. [38, 33] have some desirable properties for video modelling (constant inference cost per timestep independent of the length of the video), but they are slow to train due to their sequential nature and have difficulties in learning over long complex sequences. Transformers [42] have emerged as a very powerful family of models for all modalities, with impressive scaling capabilities. However, they have a significant memory footprint and latency due to the quadratic complexity of the self-attention operation. Recently, a new family of linear recurrent networks [17, 16, 32, 4], referred to as State Space Models (SSMs), has emerged as an answer to the quadratic complexity of self-attention and the slow training of RNNs, with promising results for vision and language  [9, 27].
@@
In this paper, we propose a hybrid architecture that combines the best of all worlds. It alternates gated linear recurrent units (LRUs) [9] applied over time, with self-attention blocks over space, and MLP over feature channels. As opposed to space and channels, time has a natural order (”arrow-of-time”) that LRUs can implicitly and efficiently model with $O(N)$ complexity in the number of input frames at training time and $O(1)$ complexity at inference time, making it possible to process videos that extend even indefinitely. Space, on the other hand, has a fixed limited dimension, for which the quadratic cost of self-attention is more accessible. From a practical perspective, using self-attention over space allows us to naturally process in parallel all the pixels of a given frame, without having to commit to a particular scanning order [27], making better use of hardware when parallel resources are available.
@@
To further limit the self-attention cost, we use spatial patches as introduced in the successful ViT [12] model. But, compared to existing video transformer models, e.g. ViViT [1], the patches do not have a fixed temporal extent. Instead, the embeddings of the spatial patches are integrated continuously into the hidden state of the gated LRUs, providing persistent memory of the entire temporal sequence up to the current frame. Furthermore, similar to convolutional networks, the parameters of the LRUs are shared over space, preventing the number of parameters from exploding as the resolution of the video increases.
@@
We refer to the resulting model as Temporal Recurrent Video Transformer (TRecViT). TRecViT is highly flexible and can address various video understanding tasks, both sparse (e.g. video classification) and dense (e.g. point tracking), trained in a supervised or self-supervised manner, e.g. using masked auto-encoding. In all our experiments, we use a causal setup that respects the arrow of time, so the model is suitable for any downstream applications, from e.g. video classification where we have offline access to the videos, to e.g. robotics, where online processing is required. Overall, our model is significantly more efficient in both memory footprint and FLOPs compared to vanilla transformers.
Paper structure: We discuss related works in more depth in section 2 and we introduce the proposed model in section 3. We discuss training regimes and analyse efficiency when comparing to baselines in section 4. In section 5, we present extensive experiments for various training regimes, different tasks and datasets. We conclude in section 6 with a discussion of the limitations of the proposed approach and directions for future work.
@@ 2 Related work
Transformers for Video. Proposed initially as language models, transformers [42] have quickly become the dominant architecture across multiple modalities (images, audio, video). Transformer blocks alternate between a spatial mixing block represented by self-attention and a (feature) channel mixing block, represented by a gated MLP. Given that the self-attention layer treats the input tokens as a set, positional encodings must be used in order to specify the location of each token. It also means that no parsing order is needed, unlike the case with RNNs. Vision transformers (ViT) [12, 29] split images into a fixed number of patches that are projected into an embedding space to obtain tokens and these are then processed by a regular transformer. Several works extended ViT to video, e.g. by replacing the regular image patches with spatio-temporal ones. The main challenge with transformers, particularly for video, is the quadratic complexity in the number of input tokens. Multiple approaches have been proposed to address this: e.g. factorisations of the self-attention operation [1, 5], iterative attention [23], sparse sampling of the input frames [34], and distributed self-attention operations across different devices [28]. Our proposed model uses a novel space-time factorisation, where the temporal dimension is handled by LRUs and the spatial dimension by self-attention.
@@
As these models scale successfully to large number of parameters, their data needs are efficiently met by using self-supervised pre-training like masked autoencoding (MAE) [40] or contrastive learning [45]. Due to the factorisation used in our architecture, using such pre-training strategies is straightforward and we include successful experiments with MAE pre-training in Section 5.
@@
SSM, a type of Linear Recurrent Model. While transformers [42] can be efficiently parallelised during training, at inference they need to pay a quadratic cost in the sequence length. On the other hand, recurrent networks [13, 20, 30, 3, 39] are compact and efficient at inference but slow at training. State Space Models (SSMs) [17, 18, 32], a particular type of linear recurrent networks, have recently been proposed as an answer to the scalability problem of RNNs, and have shown strong performance in language and other long-range dependencies tasks [9, 16].
SSMs, like S4 [18], S4D [19], or Mamba [16] have been introduced as particular discretizations of a continuous time linear system. On the other hand, the linear recurrent unit (LRU) [32] was designed by identifying the minimal set of changes to a vanilla RNN [13] that allows it to obtain the same key properties as the S4D architecture [18]; we discuss the LRU in more detail in Section 3. Improving on the LRU, the gated LRU [9] introduces gating mechanisms similar to LSTM or GRU architectures, to filter the input sequence, while the recurrent gate controls the rate of the information decay. Importantly, different from LSTM or GRU, these gates do not depend on the previous state, which would prevent parallelisation at training time. In our work, we use gated LRUs, but we expect similar results to be obtained when using other gated SSM blocks like Mamba within our factorisation.
@@
SSMs for Video. While SSMs have mostly been explored in language, several architectures like S4 and Mamba have also been adapted to image and video modalities [44]. ViS4mer [21] uses a ViT image encoder to process videos frame by frame, and integrates their representations over time using S4 blocks at the top. TranS4mer [22] uses self-attention over short clips and integrates these with gated S4 blocks. More recently, the Mamba architecture was extended to images and videos by having it process a flattened 1D sequence of image or video patches. This requires defining a processing order for the patches, and different orders have been proposed, e.g. bidirectional and following a column or row order [46, 27]. As opposed to these Mamba-based architectures, our factorisation naturally uses the arrow-of-time to decide the scanning order, resulting in a causal model. Another important benefit of our hybrid architecture is that we can initialise the ViT blocks with strong existing pre-trained weights. This leads to strong performance even at larger scale, as opposed to VideoMamba [27] where the authors report severe overfitting issues, requiring distillation from smaller models when training in a supervised fashion or distillation from CLIP features [37] for self-supervised training.
@@ 3 TRecViT Architecture
The proposed architecture, TRecViT, is composed of repeated identical blocks, each performing a sequence of information mixing steps across the different dimensions of the video signal: time, space, and channels; see Figure 1. The mixing over the time dimension is handled by gated linear recurrent units (LRUs), similar to the one introduced in [9] for language. Each spatial token is associated with an LRU that processes the tokens within the same temporal tube over time, without mixing the information across temporal tubes. The LRUs share parameters over space, similar to a convolutional network. When applying this temporal mixing operation, the space dimension is transposed into the batch dimension. The mixing over spatial and channel dimensions is handled by a standard ViT block, which first performs the spatial mixing through a self-attention operation, then the channel mixing by using an MLP. When performing the spatial and channel mixing, the time dimension is transposed into the batch dimension.
@@
Empirically, we show that this factorization and choice of building blocks is more efficient for understanding temporal dynamics compared to video transformer approaches (e.g. ViViT [1]) or pure SSM models. By applying self-attention over the spatial dimensions, we allow all tokens to attend to all the other tokens in parallel, without having to commit to a particular order (unlike in VideoMamba). We employ strong transformer blocks from ViTs for this operation, including their Imagenet pre-trained weights. The recurrence of the temporal processing enables efficient frame-by-frame inference over long videos, with constant memory footprint and causal operation.
3.1 Background on LRUs
Linear Recurrent Units (LRUs) [32], similar to SSMs, belong to the family of linear recurrent architectures and have been shown to be competitive with Transformers on language tasks [9, 16]. One potential interpretation of the success of these models, as outlined in [32], is that by sacrificing the nonlinear recurrence typical of a recurrent model, we can improve the scalability and controllability of the system. Specifically, the linearity allows the recurrent matrix to be diagonalised through eigenvalue decomposition and absorbing the (dense) eigenvectors matrix into the neighbouring layers. This gives direct access to the eigenvalues of the Jacobian of the transfer function characterising the system. By initialising these eigenvalues within the unit circle, we have guaranteed stability of the system, bypassing issues like vanishing or exploding gradients. In addition, through the specific initialisation range of the eigenvalues within [0, 1] we can control how quickly the information vanishes, with eigenvalues close to 1 promoting longer-term memory. However, using only linear recurrence can greatly limit the expressivity of the layer. In  [31], the authors show that by using these layers within a typical transformer structure that alternates linear recurrences with point-wise nonlinearities (e.g. the MLP block), the overall architecture can be shown to be a universal approximator of finite sequence-to-sequence maps.

### v2

@@ Abstract
We propose a novel block for causal video modelling. It relies on a time–space–channel factorisation with dedicated blocks for each dimension: gated linear recurrent units (LRUs) perform information mixing over time, self-attention layers perform mixing over space, and MLPs over channels. The resulting architecture TRecViT is causal and shows strong performance on sparse and dense tasks, trained in supervised or self-supervised regimes, being the first causal video model in the state-space models family. Notably, our model outperforms or is on par with the popular (non-causal) ViViT-L model on large scale video datasets (SSv2, Kinetics400), while having $3\times$ less parameters, $12\times$ smaller memory footprint, and $5\times$ lower FLOPs count than the full self-attention ViViT, with an inference throughput of about 300 frames per second, running comfortably in real-time. When compared with causal transformer-based models (TSM, RViT) and other recurrent models like LSTM, TRecViT obtains state-of-the-art results on the challenging SSv2 dataset. Code and checkpoints are available online https://github.com/google-deepmind/trecvit.
@@ 1 Introduction
Video understanding requires low-level scene understanding (e.g. how objects move) and high-level reasoning (e.g. causal relations between events) over a signal that is high-dimensional, can be noisy, and contains high correlations and redundancies in both spatial and temporal dimensions. Efficient video modelling needs high-capacity models that can represent the sheer diversity and richness of real-world videos, while operating in a causal manner with reasonable compute and memory footprint both at training and during inference time – these efficiency requirements are critical for Robotics or augmented reality applications. Convolutional neural networks (Carreira and Zisserman, 2017b; Feichtenhofer et al., 2019; Lin et al., 2019; Kwon et al., 2020) have been a successful family of causal models for video, but their scaling capabilities (in both data and parameters) are limited due to their inductive biases (locality, invariance). Recurrent neural networks, e.g. (Srivastava et al., 2015; Patraucean et al., 2016) have some desirable properties for video modelling (constant inference cost per timestep independent of the length of the video, causality for unidirectional models), but they are slow to train due to their sequential nature and have difficulties in learning over long complex sequences. Transformers (Vaswani et al., 2017) have emerged as a very powerful family of models for all modalities, with impressive scaling capabilities. However, they have a significant memory footprint and latency due to the quadratic complexity of the self-attention operation and their performance degrades when using causal self-attention masks. Recently, a new family of linear recurrent networks (Gu et al., 2020; Gu and Dao, 2023; Orvieto et al., 2023b; Beck et al., 2024), referred to as State Space Models (SSMs), has emerged as an answer to the quadratic complexity of self-attention and the slow training of RNNs, with promising results for vision and language  (De et al., 2024; Li et al., 2024). However, none of the existing video SSM architectures can run in a causal manner, their performance strongly depends on bidirectional operation.
@@
In this paper, we propose a hybrid architecture that combines the best of all worlds. It alternates gated linear recurrent units (LRUs) (De et al., 2024) applied over time, with self-attention blocks over space, and MLP over feature channels. As opposed to space and channels, time has a natural order ("arrow-of-time") that LRUs can implicitly and efficiently model in a causal manner with $O(N)$ complexity in the number of input frames at training time and $O(1)$ complexity at inference time, making it possible to process in real-time videos that extend even indefinitely. Space, on the other hand, has a fixed limited dimension, for which the quadratic cost of self-attention is more accessible. From a practical perspective, using self-attention over space allows us to naturally process in parallel all the pixels of a given frame, without having to commit to a particular scanning order (Li et al., 2024), making better use of hardware when parallel resources are available. Importantly, by restricting the LRUs to temporal-only recurrence, this factorisation reduces the sequence length by about two orders of magnitude compared to models that apply recurrence across both space and time (Zhu et al., 2024; Li et al., 2024). Such models typically require bidirectional scanning to attain strong performance, preventing them from operating in a causal manner.
@@
To further limit the self-attention cost, we use spatial patches as introduced in the successful ViT (Dosovitskiy et al., 2021) model. But, compared to existing video transformer models, e.g. ViViT (Arnab et al., 2021), the patches do not have a fixed temporal extent. Instead, the embeddings of the spatial patches are integrated continuously into the hidden state of the gated LRUs, providing persistent memory of the entire temporal sequence up to the current frame, leading to a causal architecture. Furthermore, similar to convolutional networks, the parameters of the LRUs are shared over space, preventing the number of parameters from exploding as the resolution of the video increases.
@@
We refer to the resulting model as Temporal Recurrent Video Transformer (TRecViT). TRecViT is highly flexible and can address various video understanding tasks, both sparse (e.g. video classification) and dense (e.g. point tracking), trained in a supervised or self-supervised manner, e.g. using masked auto-encoding. In all our experiments, we use a causal setup that respects the arrow of time, so the model is suitable for any downstream applications, from e.g. video classification where we have offline access to the videos, to e.g. Robotics, where online processing is required. Overall, our model is significantly more efficient in both memory footprint and FLOPs compared to vanilla transformers, and obtains state-of-the-art performance when compared to causal models, while comfortably running in real time (e.g. throughput of about 300 frames per second for point tracking task).
Contributions: We propose a causal video architecture that uses a novel factorisation interleaving LRUs and ViT blocks. We run ablations for the building blocks to find optimal hyperparameters and present extensive experiments on multiple tasks and training regimes, showing the versatility and strong performance of our model.
@@ 2 Related work
Transformers for Video. Proposed initially as language models, transformers (Vaswani et al., 2017) have quickly become the dominant architecture across multiple modalities (images, audio, video). Transformer blocks alternate between a spatial mixing block represented by self-attention and a (feature) channel mixing block, typically represented by a gated MLP. Given that the self-attention layer treats the input tokens as a set, positional encodings must be used in order to specify the location of each token. This also implies that no parsing order is needed, unlike the case with RNNs. Vision transformers (ViT) (Dosovitskiy et al., 2021; Liu et al., 2021) split images into a fixed number of patches that are projected into an embedding space to obtain tokens and these are then processed by a regular transformer. Several works extended ViT to video (Arnab et al., 2021; Bertasius et al., 2021; Li et al., 2021; Fan et al., 2021; Patrick et al., 2021), e.g. by replacing the regular image patches with spatio-temporal ones. The main challenge with transformers, particularly for video, is the quadratic complexity in the number of input tokens. Multiple approaches have been proposed to address this: e.g. factorisations of the self-attention operation (Arnab et al., 2021; Bertasius et al., 2021), iterative attention (Jaegle et al., 2022), sparse sampling of the input frames (Piergiovanni et al., 2023), and distributed self-attention operations across different devices (Liu et al., 2024). Our proposed model uses a novel space-time factorisation, where the temporal dimension is handled by LRUs and the spatial dimension by self-attention, resulting in a causal efficient video architecture.
@@
As these models scale successfully to large number of parameters, their data needs are efficiently met by using self-supervised pre-training like masked autoencoding (MAE) (Tong et al., 2022) or contrastive learning (Zhao et al., 2024). Due to the factorisation used in our architecture, using such pre-training strategies is straightforward and we include successful experiments with MAE pre-training in Section 5.
@@
SSM, a type of Linear Recurrent Model. While transformers (Vaswani et al., 2017) can be efficiently parallelised during training, at inference they need to pay a quadratic cost in the sequence length. On the other hand, recurrent networks (Elman, 1990; Hochreiter and Schmidhuber, 1997a; Mikolov et al., 2010; Bahdanau et al., 2014; Sutskever et al., 2014) are compact and efficient at inference but slow at training. State Space Models (SSMs) (Gu et al., 2020; 2021; Orvieto et al., 2023b), a particular type of linear recurrent networks, have recently been proposed as an answer to the scalability problem of RNNs, and have shown strong performance in language and other long-range dependencies tasks (De et al., 2024; Gu and Dao, 2023).
SSMs, like S4 (Gu et al., 2021), S4D (Gu et al., 2022), or Mamba (Gu and Dao, 2023) have been introduced as particular discretizations of a continuous time linear system. On the other hand, the linear recurrent unit (LRU) (Orvieto et al., 2023b) was designed by identifying the minimal set of changes to a vanilla RNN (Elman, 1990) that allows it to obtain the same key properties as the S4D architecture (Gu et al., 2021). Specifically, the nonlinear recurrence typical of a recurrent model was removed to improve the scalability and controllability of the system, as the linearity allows the recurrent matrix to be diagonalised through eigenvalue decomposition and absorbing the (dense) eigenvectors matrix into the neighbouring layers. This gives direct access to the eigenvalues of the Jacobian of the transfer function characterising the system. By parametrising the diagonal weight matrix containing the eigenvalues such that all entries are constrained to be below one, the system is guaranteed to be stable, bypassing exploding gradients issues. However, using only linear recurrence can greatly limit the expressivity of the layer. In  (Orvieto et al., 2023a), the authors show that by using these layers within a typical transformer structure that alternates linear recurrences with point-wise nonlinearities (e.g. the MLP block), the overall architecture recovers expressivity through depth and can be shown to be a universal approximator of finite sequence-to-sequence maps.
Improving on the LRU, the gated LRU (De et al., 2024) introduces gating mechanisms similar to LSTM or GRU architectures, to filter the input sequence, or, for the recurrent gate, to control the rate of the information decay. Importantly, different from LSTM or GRU, these gates do not depend on the previous state, which would prevent parallelisation at training time. A series of recent works rely on formulating the linear recurrence as a form of linear attention, such as Mamba 2 (Dao and Gu, 2024), Gated Delta Networks (Yang et al., 2025), MesaNetworks (von Oswald et al., 2025), akin to previous observation, providing a more direct connection to attention. These works exploit this connection to show that the recurrence can be seen as an update rule of a local objective, similar to interpretability works on in-context learning. In this line of work, it is also worth noting xLSTMs (Beck et al., 2024), which connects SSMs more explicitly to traditional non-linear models such as LSTMs (Hochreiter and Schmidhuber, 1997b). In our work, we use gated LRUs, but we expect similar results when using other gated SSM blocks like Mamba or xLSTM within our factorisation.
@@
SSMs for Video. While SSMs have mostly been explored in language, several architectures like S4 and Mamba have also been adapted to image and video understanding (Zhang et al., 2024) and generation (Gao et al., 2024). ViS4mer (Islam and Bertasius, 2022) uses a ViT image encoder to process videos frame by frame, and integrates their representations over time using S4 blocks at the top. TranS4mer (Islam et al., 2023) uses self-attention over short clips and integrates these with gated S4 blocks. More recently, the Mamba architecture was extended to images and videos by having it process a flattened 1D sequence of image or video patches. This requires defining a processing order for the patches, and different orders have been proposed, e.g. bidirectional and following a column or row order (Zhu et al., 2024; Li et al., 2024; Shi et al., 2025; Lu et al., 2025). As opposed to these Mamba-based architectures, our factorisation naturally uses the arrow-of-time to decide the scanning order, resulting in a causal model. Another important benefit of our hybrid architecture is that we can initialise the ViT blocks with strong existing pre-trained weights. This leads to strong performance even at larger scale, as opposed to VideoMamba (Li et al., 2024) where the authors report severe overfitting issues, requiring distillation from smaller models when training in a supervised fashion or distillation from CLIP features (Radford et al., 2021) for self-supervised training.
Causal video models. All the models mentioned above process the videos in an offline manner, i.e. require access to all frames at the same time, making their application in streaming setups (e.g. Robotics, virtual reality) sub-optimal. A few works have proposed adaptations of convnets and Transformers to causal operation (Lin et al., 2019; Zhao et al., 2023; Yang et al., 2022). In (Zhao et al., 2023), the authors store the keys and values from previous frames to perform cross attention with queries from the current frame, but they still use a non-causal temporal transformer for action recognition tasks. The closest to our work is RViT (Yang et al., 2022), which proposes a linear attention-based gating for processing videos in a recurrent manner. In the experiments section, we show that our recurrence based on LRUs leads to state-of-the-art results in causal operation settings and competitive performance with non-causal models.
@@ 3 TRecViT Architecture
Figure 1: Left: TRecViT architecture. Each video frame is divided into non-overlapping patches that are linearly projected into a token embedding space. We then add a learnt spatial positional encoding. The tokens are passed through gated linear recurrent units (LRUs) that share parameters across space. The outputs of the recurrent blocks are then processed by a ViT block. The recurrent operation followed by ViT is repeated N times. Right: TRecViT block. The input is a batch of videos, each frame with N tokens. We apply recurrent units over temporal tubes to integrate information over time, and self-attention and MLP across tokens within each frame. Note that the recurrent units share parameters, but the information is not mixed across temporal tubes. Similarly, the ViT blocks share parameters, but the information is not mixed across frames. There are skip connections around each block, not included here to reduce clutter.
Notations: Let $X\in[0,1]^{T\times H\times W\times 3}$ be an RGB video with $T$ frames and $H\times W$ pixels. The video frames are split into $N$ non-overlapping patches $p_{t}^{k}$ of size $n\times n\times 3$, with $t\in\{1,T\}$ and $k\in\{1,N\}$. Let $x_{t}^{k}$ be the tokens obtained after the linear projection of the patches and the addition of the spatial positional encoding, with token size $1\times 1\times d$, where $d$ is the token feature dimension. A temporal tube is a sequence $\{x_{t}^{k}|t=\overline{1,T}\}$ containing the tokens from the same spatial location across frames.
@@
The proposed architecture, TRecViT, is composed of repeated identical blocks, each performing a sequence of information mixing steps across the different dimensions of the video signal: time, space, and channels; see Figure 1. The mixing over the time dimension is handled by gated linear recurrent units (LRUs), similar to the one introduced in (De et al., 2024) for language. Each spatial token is associated with an LRU that processes the tokens within the same temporal tube over time, without mixing the information across temporal tubes, so each LRU has its own state – we use LRU1, LRU2,… in Figure 1 to highlight this. The LRUs share parameters over space, similar to a convolutional network. When applying this temporal mixing operation, the space dimension is transposed into the batch dimension.
The mixing over spatial and channel dimensions is handled by a standard ViT block, which first performs the spatial mixing through a self-attention operation, then the channel mixing by using an MLP. When performing the spatial and channel mixing, the time dimension is transposed into the batch dimension. We first perform temporal mixing, followed by spatial and channel mixing. We found this time-space order to produce better results compared to space-time order. We hypothesise that, by applying LRUs first, this allows them to focus on more local, easier to model, information at the first layer, instead of operating directly on features that mix information across the entire frame.
@@
Empirically, we show that this time-space-channel factorization and choice of building blocks is more efficient for understanding temporal dynamics compared to video transformer approaches (e.g. ViViT (Arnab et al., 2021)) or pure SSM models. By applying self-attention over the spatial dimensions, we allow all tokens to attend to all the other tokens in parallel, without having to commit to a particular order (unlike in VideoMamba). We employ strong transformer blocks from ViTs for this operation, including their Imagenet pre-trained weights. The recurrence of the temporal processing enables efficient frame-by-frame inference over long videos, with constant memory footprint and causal operation.